OpenStudy (anonymous):

Make a box and whisker plot. Dog weights(lb): 23,15,88,34,33,49,52,67,42,71,28

OpenStudy (anonymous):

Please and thank you

OpenStudy (the_fizicx99):

Find the maximum, minimum, Q1, and Q3 then plot it, find the IQR if you'd like, it's Q3 - Q1.

OpenStudy (the_fizicx99):

Do you know how to find Q1 and Q3?

OpenStudy (anonymous):

no do i have to put then in least to greatest?

OpenStudy (the_fizicx99):

Yes, that is very helpful. Can you put them in least to greatest?

OpenStudy (anonymous):

So, 15, 23,28,33,34,42,49,71,88

OpenStudy (the_fizicx99):

I'm going to trust that you put them correctly xD There's 9 so it's odd, in your favor, Grab the first 4 terms, and leave the middle one alone, then grab the last 4. {15, 23,28,33} Middle: 34 {42,49,71,88} Now find the median of {15, 23,28,33} and {42,49,71,88} :) The median is the middle term, I'll do {15, 23,28,33}, because you have two numbers in the middle, you have to take 23 and 28 and add them, 23 + 28 = 51/2 = 25.5 Now you do the other! {42,49,71,88}

OpenStudy (anonymous):

I MEST UP UGHHHH!:0 i forgot some

OpenStudy (the_fizicx99):

:O

OpenStudy (anonymous):

i kno ugh

OpenStudy (anonymous):

15,23,28,33,34,42,49,52,71,88

OpenStudy (anonymous):

sorry i missed two. ugh my head hurts i been doing math all day!

OpenStudy (the_fizicx99):

Crap :O Well, now it's even xD same process, grab 5, {42,49,52,71,88} and {15,23,28,33,34} Find the median of them, in a way it's easier, the first one is 52 and the last one is?

OpenStudy (anonymous):

28?

OpenStudy (the_fizicx99):

Yesh ;) So, Q1 is 28 and Q3 is 52, Draw a horizontal rectangle above a number line that has a scale to fit your data. The left end of the rectangle = Q1. The right end = Q3. Draw a vertical line inside the rectangle where the median value occurs.

OpenStudy (anonymous):

can u draw it for me? please

OpenStudy (the_fizicx99):

I'm a horrible artist ;_; it would look really bad, it'd be very messy :(

OpenStudy (anonymous):

its ok at least an idea please

OpenStudy (the_fizicx99):

http://prntscr.com/3cgv5x There's an example, fairly easy :>

OpenStudy (anonymous):

what numbers do u prefer on my number lines?

OpenStudy (the_fizicx99):

What do you mean :o

OpenStudy (anonymous):

on my number line what do i number them as?

OpenStudy (anonymous):

idk. can u just draw the whisker plot for me please?:)

OpenStudy (the_fizicx99):

1. Find the MAX (highest number on your data), then find the MIN (lowest point in your data), draw a horizontal line starting from the min up to the max, draw a boxish like rectangle on the median of the data (Q2) 2. Label the Q1 and Q3 on your box and that's it :)

OpenStudy (anonymous):

? hum

OpenStudy (the_fizicx99):

 Something like that xD

OpenStudy (anonymous):

do i have to put numbers on the line? idk ugh

OpenStudy (the_fizicx99):

Yesh go on from 0 - 100 at intervals of 5 0 - 5 - 10 - 15 - 20 - 25 etc.